# Exploring Concept Depth: How Large Language Models Acquire Knowledge at Different Layers?

### v1

@@ Abstract
This paper studies the phenomenon that different concepts are learned in different layers of large language models, i.e. more difficult concepts are fully acquired with deeper layers. We define the difficulty of concepts by the level of abstraction, and here it is crudely categorized by factual, emotional, and inferential. Each category contains a spectrum of tasks, arranged from simple to complex. For example, within the factual dimension, tasks range from lie detection to categorizing mathematical problems. We employ a probing technique to extract representations from different layers of the model and apply these to classification tasks. Our findings reveal that models tend to efficiently classify simpler tasks, indicating that these concepts are learned in shallower layers. Conversely, more complex tasks may only be discernible at deeper layers, if at all. This paper explores the implications of these findings for our understanding of model learning processes and internal representations. Our implementation is available at https://github.com/Luckfort/CD.
[…]
123
@@ 1 Introduction
Modern large language models (LLMs) such as GPT-4 [1], Gemma [29], Llama [32] have impressive generation and inference capabilities. It is a common insight that large models encode a large amount of knowledge within their parameters and that a larger number of parameters leads to better performance. Consequently, popular large models are usually very deep networks containing billions or even trillions of parameters [16, 13]. It is observed that as the size of the model parameters increases, the model exhibits dramatic and unpredictable performance improvements on specific tasks, which are so-called emergent abilities [33]. For instance, GPT-3 [6] shows a large increase in performance after exceeding 13B parameter in size, and a similar phenomenon was observed for LaMDA [31] after exceeding 68B parameter [34]. However, it remains obscure how and where LLMs actually understand the concept of knowledge correctly, and there is currently a lack of empirically summarized research on the relationship between LLMs’ network depth and conceptual understanding.
Recent studies usually approach concerns above from the following two perspectives, 1) analyzing model weights and architectures [18], e.g., the pruning method removes weights based on their importance and reverse inference of where the model learns the concept of knowledge by determining whether it will affect the model performance. No matter it is single pruning [18] or iterative pruning [15], existing pruning methods applied to LLMs are usually designed to be relatively complex and not generalizable, requiring adjustments for different datasets [23]. 2) probing representations [2], which typically utilize linear classifier probes to test the performance of the hidden layer representations. For example, the most recent probing-based work [16] explores each layer’s performance of the Llama series on the counterfactual dataset [24] with representations from different layers.
In this paper, we propose the notion of "Concept Depth" to measure where different models, on different types of datasets, understand knowledge. To our knowledge, this is the first time this notion has been introduced in related work. We draw empirical summaries from extensive experiments on the "Concept Depth" of existing mainstream large language models on different types of datasets through probe techniques. As shown in Figure 1, we capture the feature responses of different layers of the LLMs for different datasets and use independent linear probes to indicate the best performance that the current layer can achieve. The core contribution of our work can be summarized as follows:
@@
Figure 1: a) The overview of process of our analysis. LLMs respond to text prompts. The probing process evaluates the optimal performance that the current layer of LLMs can achieve. b) A conceptual example of layer-wise representations for probe detection. Orange points and Blue points represent fake samples and real samples, respectively. The probing task here is the binary classification on different types of samples.
Concept Depth. Introducing the concept of "concept depth" to measure where a concept is learned by different size LLMs. Our experiments show that basic concepts are often learned at a low level of "concept depth", while more complex concepts require more depth. This is a consistent phenomenon across LLMs of different model families and different sizes.
Deconstructing LLMs Capabilities. Deconstructing LLM capabilities by analogy with the composition of complex human capabilities. We conduct experiments on a large number of relevant datasets, synthesize comprehensive performance summaries of different LLMs, and obtain specific concept depths of optimal performance for different capabilities.
Robustness from Concept Depth perspective. Providing a new perspective on LLMs robustness. We conduct ablation experiments on other factors (e.g., quantization and random noise) that may affect the accuracy of LLMs inference, which showed that these interfering factors affect the concept depth of the model so that the LLMs appear to end up learning concepts at a deeper location than before. It can be assumed that deeper models have more concept depth to counteract interference.
@@ 2.1 Concepts representation in DNNs
Identifying similarities across various examples to form concepts, plays a crucial role in human decision-making [3]. Many studies have explained DNNs’ (Deep Neural Networks) decision-making based on a conceptual perspective, describing the global behavior of DNNs in a human-understandable way [35, 27, 28, 8]. For example, Yeh et al. [35] demonstrated that DNNs exhibit conceptual representations through the activation patterns observed in their hidden or output layers. Further, Räz [27] indicates that DNNs learned not only conceptual representations of predicted categories but also indirect concepts that contribute to the prediction. A notable study reveals the existence of a ‘representation bottleneck,’ highlighting a cognitive disparity between DNNs and humans. This phenomenon is characterized by DNNs’ tendency to grasp representations of concepts that are either too simple or overly complex, while they often struggle with acquiring representations of concepts of moderate complexity [8]. Motivated by previous work, our paper aims to cover concepts of different complexities to further understand concepts within LLMs.
@@ 2.2 Knowledge and concepts in LLMs
The great performance of the LLMs in various downstream tasks (e.g. LLMs can predict factual statements about the world based on prompts [24].) has led to a great discussion about whether these capabilities are ‘stochastic parrots’ or LLMs really understands these concepts. Pioneeringly, Gurnee and Tegmark [14] showed that LLMs internally store concepts like latitude, longitude, and time. Similarly, another work showed that the truth of a statement can be detected from the internal state of LLMs [4]. Geva et al. [12] also came to similar conclusions by artificially blocking or “knocking out" specific parts of the LLMs to observe their effects on the inference process. These related studies show the existence of structures for understanding concepts within LLMs, which motivates us to explore how concepts at various complexities are encoded within various depths of LLMs.
@@ 2.3 Explorations of Interpretability in LLMs
A large number of related studies have deconstructed the inner layers of LLMs from various perspectives to understand the mechanisms inside such models [36]. Fan et al. [10] computes stopping signals by evaluating key features to early stop the LLM inference and get the internal performance of the LLMs, concluding that not all layers are necessary. Through pruning the LLMs, Gromov et al. [13] found that the parameters of some layers were not utilized correctly. Men et al. [23] also shows a high level of redundancy in the LLM architecture. Probes trained with logistic regression are a well-established method [2] that has been applied in classifying the truthfulness of LLMs and has been validated in many studies [22, 4, 19]. The latest work related to us detects different layers in Llama series responding to facts or counterfactuals by probing techniques [16]. Inspired by these work, we propose Concept Depth to summarize these phenomena. Our work is different from previous work. Our work is relatively comprehensive, and we focus on the Concept Depth of where it appears in the LLMs, analyzing it experimentally by training linear classifier probes.
Figure 2: The LLMs are trying to understand easy tasks and complex tasks. The more complex the task is, the deeper layers an LLM needs to understand it. The stronger the LLM is, the more complex the task level it can learn.
@@ 3 Analyzing Method
In this paper, we designed a complete probing framework to under-stand how concepts at various levels are encoded within LLMs andinvestigate whether LLMs’ internal representations are robust withrespect to concept.
3.1 Visualizing Layer-wise Representations
Contemporary LLMs typically consist of numerous layers of deep Transformer decoders (see Table 1), yet the conceptual significance of the encoded information within each layer remains largely unexplored. To illustrate this point, we analyzed the representations from every layer across an entire LLM, employing principal component analysis (PCA) to facilitate visualization. Specifically, for each layer’s representation, we will downscale the representation to two dimensions and visualize it following the gold label (e.g., math problem or not). The comparison across layers at various depths will indicate LLMs’ understanding of problems in need of different concepts. For instance, as shown in Figure 1, we project the representation of Counterfact [24] in the 80% depth of Gemma-7b, and the label is the real answer of the data set. It can be observed that gemma-7b has significantly different responses for inputs with different type answers at this depth.
@@ 3.2 Linear Classifier Probing
To better understand how representations from different layers could contribute to the final prediction, we utilized a linear classifier to further quantify the probing differences. For one specific task $w$, the hidden feature set in LLMs is $x\in\mathbb{R}^{n\times d_{model}}$, where $n$ denotes number of samples, and $x^{(i)}\in\mathbb{R}^{1\times d_{model}}$ represent the representation at a certain layer, where $d_{model}$ donate the dimension for the hidden layer representation and the details of $d_{model}$ for each LLM can be found in the appendix. Binary label $y^{(i)}$ is set as 0 or 1. The objective function of such a binary Logistic regression classifier probes with L2 regularization can be written as:
$\scriptsize J(\theta)=-\frac{1}{n}\sum_{i=1}^{n}Cost(\sigma(x^{(i)}),y^{(i)})+\frac{\lambda}{2n}\sum_{j=1}^{m}\theta_{j}^{2}$ (1)
and
$\scriptsize Cost(\sigma(x^{(i)}),y^{(i)})=y^{(i)}\log\left(\sigma(\theta^{T}x^{(i)})\right)+(1-y^{(i)})\log\left(1-\sigma(\theta^{T}x^{(i)})\right)$ (2)
where $\theta$ is the parameter in this Logistic regression model, $\lambda$ is the regularization parameter. The linear model predicts LLM’s response to the test set, compared with the true label. This yields a quantification of how well LLMs understand the current depth.
Refer to caption (a) Refer to caption (b)
Refer to caption (c) Refer to caption (d)
Figure 3: Analysis diagrams of Section 5.2. The converge proportion and peak accuracy of each model over the nine datasets. (a) and (c) shows the converge proportion over the datasets. (b) and (d) shows the peak accuracy over the datasets.
@@ 4 Experimental Setting
Our experiments used 9 datasets containing three aspects (human emotions, logical reasoning, and factual analysis) and categorized these 9 datasets into easy and complex datasets according to the performance of the LLMs on each dataset. Specifically, datasets in which the linear probes can obtain high classification accuracy at the initial or middle depth of the LLMs are categorized as easy datasets while other datasets with large fluctuations and stable classification accuracy occurring at the deep depth of the LLMs are categorized as complex datasets. In Section 4.1, we introduce the LLMs used for experiments. And give a description of these nine datasets in Section 4.2.
@@ 4.1 Models
In this paper, we employ three families of open-source models: Gemma (2B, 7B) [30], LLaMA (7B, 13B) [32], and QWen (0.5B, 1.8B, 4B, 7B, and 14B) [5] to support our analysis. The number of layers in each model is shown in Table 1. During probing analysis, we choose a linear classifier probe in the experiments. The ratio of training set to test set is 8:2. We follow the usual approach of LLMs probe classifier [9, 26], where the feature representation of the final transformer layer at each layer of LLMs(e.g. 14-th ’post_attention_layernorm’ in Llama2-7b (32 Layers in total) is as input to the probe. We provide the detailed structure of Llama2-7b model in appendix. The other series of models follow a similar structure.
@@
Gemma-2B 18 Gemma-7B 28 LLaMA-7B 32
LLaMA-13B 40 QWen-0.5B 24 QWen-1.8B 24
QWen-4B 40 QWen-7B 32 QWen-14B 40
@@ 4.2 Datasets
We used nine datasets on fact (Cities[22], CommonClaim[7], Counterfact[24]), emotion (STSA[17], IMDb[20], Sarcasm[25], HateEval [21]), and inference (StrategyQA[11], Coinflip[34]) for our experiments, and sample examples of each are given in the Appendix, Section A.2, following is a brief description of these datasets. The number of samples we use for each dataset is shown at Table 2.
Table 2: Number of samples we use in each dataset.
Dataset Size Dataset Size Dataset Size
Cities 1496 CommonClaim 6000 Counterfact 4000
HateEval 6000 STSA 6920 IMDb 2000
Sarcasm 6000 StrategyQA 2290 Coinflip 500
Cities [22]: consists of statements about the location of cities and their veracity labels (e.g., The city of Zagreb is in Japan, which is wrong).
CommonClaim [7]: A dataset of boolean statements, each labeled by two humans as common-knowledge-true, common-knowledge-false, or neither.
Counterfact [24]: Counterfact includes thousands of counterfactuals along with text that allows quantitative testing of specificity and generalization when learning a counterfactual.

### v4

@@ Abstract
Large language models (LLMs) have shown remarkable performances across a wide range of tasks. However, the mechanisms by which these models encode tasks of varying complexities remain poorly understood. In this paper, we explore the hypothesis that LLMs process concepts of varying complexities in different layers, introducing the idea of “Concept Depth” to suggest that more complex concepts are typically acquired in deeper layers. Specifically, we categorize concepts based on their level of abstraction, defining them in the order of increasing complexity within factual, emotional, and inferential tasks. We conduct extensive probing experiments using layer-wise representations across various LLM families (Gemma, LLaMA, Qwen) on various datasets spanning the three domains of tasks. Our findings reveal that models could efficiently conduct probing for simpler tasks in shallow layers, and more complex tasks typically necessitate deeper layers for accurate understanding. Additionally, we examine how external factors, such as adding noise to the input and quantizing the model weights, might affect layer-wise representations. Our findings suggest that these factors can impede the development of a conceptual understanding of LLMs until deeper layers are explored. We hope that our proposed concept and experimental insights will enhance the understanding of the mechanisms underlying LLMs. Our codes are available at https://github.com/Luckfort/CD.
Exploring Concept Depth: How Large Language Models Acquire Knowledge at Different Layers?
Mingyu Jin1, Qinkai Yu2, Jingyuan Huang1, Qingcheng Zeng3, Zhenting Wang1, Wenyue Hua1, Haiyan Zhao4, Kai Mei1, Yanda Meng5, Kaize Ding3, Fan Yang6, Mengnan Du4, Yongfeng Zhang1 1Rutgers University, 2University of Liverpool, 3Northwestern University, 4New Jersey Institute of Technology, 5University of Exeter, 6Wake Forest University {mingyu.jin, chy.huang, yongfeng.zhang}@rutgers.edu , sgqyu9@liverpool.ac.uk , mengnan.du@njit.edu
footnotetext: **Mingyu, Qinkai, Jingyuan, and Qingcheng are the main contributors. The order of the authors is determined by flipping a coin.**
@@ 1 Introduction
LLMs such as GPT-4 Achiam et al. (2023) and LLaMA Touvron et al. (2023) have impressive generation and reasoning capabilities Chang et al. (2023); Su et al. (2024a, c). It is widely accepted that these models embed substantial knowledge in their parameters, with performance improving as the number of parameters increases Ju et al. (2024), also known as emergent abilities Wei et al. (2022a). For instance, GPT-3 Brown et al. (2020) shows a large increase in performance after scaling up to 13B parameters, and a similar phenomenon was also observed for LaMDA Thoppilan et al. (2022) after exceeding 68B parameters Wei et al. (2022b). However, it is not well understood how LLMs accurately grasp the concept of knowledge. In this paper, we investigate the following research question: Can shallow layers in LLMs capture meaningful features of simple knowledge, while complex concepts need deeper layers to capture their meaningful features? like Figure 2. We hope to explore the connection between the depth of language models’ neural networks and their conceptual understanding ability by studying this question.
@@
Figure 1: The left figure provides an overview of our analysis process. LLMs respond to text prompts, and the probing process assesses the optimal performance achievable by the current LLM layer. The right figure illustrates a demonstration of layer-wise representations by probe detection. In this demonstration, orange points represent fake samples, while blue points represent real samples. In this case, the probe tries to classify between two categories.
Recent studies on understanding the reasoning abilities of LLMs focus on two main strategies: probing representations and model pruning. Probing involves using linear classifier probes to analyze the performance of hidden layer representations; for instance, Duan et al. (2024) examines changes in LLMs’ internal representations during hallucinations, while Ju et al. (2024) investigates the performance of different layers in the LLaMA series using synthetic counterfactual datasets. On the other hand, model pruning removes redundant parameters based on their importance in seeing if performance is significantly affected. This method, although effective, can be complex and time-consuming. For example, Zhang et al. (2023) uses gradient information to decide the pruning components, and Gromov et al. (2024) even requires QLoRA fine-tuning Dettmers et al. (2024) to do the pruning. Given these complexities, Our work primarily analyzes the representations obtained through probing techniques. Building upon previous work, we aim to gain a more comprehensive understanding of the layer representations within LLMs.
Our general framework is shown in Figure 1. We trained independent linear probes for each layer of LLMs to predict the binary label, thereby determining the optimal performance achievable with the representations of each respective layer.
Drawing from our empirical findings, we propose the notion of “Concept Depth” as a novel metric to evaluate the capability of different models in comprehending varying levels of knowledge across their layers. This is the first time such a concept has been introduced in the relevant literature.
Our empirical results ranging from 3 popular LLMs families (Gemma Team et al. (2023), LLaMA Touvron et al. (2023), and Qwen Bai et al. (2023)) and 9 datasets reveal that “Concept Depth” is widely applicable in existing mainstream LLMs. Besides, we conducted comprehensive robustness analyses, introducing random strings as the noise or quantization, to further understand how the reasoning of LLMs is sensitive to noise. To conclude, our main contributions could be summarized as follows:
Concept Depth. We introduce the idea of “Concept Depth” to measure different layers’ abilities to learn different levels of concepts. We first anchored the difficulty of the dataset using LLaMA-3-8b-Instruct Dubey et al. (2024) and then tested “Concept Depth” with other models. Our results show that simpler concepts are often learned at shallower levels, while complex concepts require deeper levels to understand like Demo 2. This phenomenon has been observed across LLMs of different model families and different sizes.
Experiments on understanding capabilities of LLMs. We experimented with multidimensional datasets (fact, emotion, and reasoning) to analyze variations in the conceptual depth of LLMs. We observed these differences across various datasets, model parameter counts, and model families (Gemma Team et al. (2023), LLaMA Touvron et al. (2023), and Qwen Bai et al. (2023)), providing a concise understanding of their impact on LLM performance and comprehension.
Robustness from Concept Depth perspective. We provide a new perspective on LLMs robustness. We conduct ablation experiments on model weight quantization and add random noise to the input that may affect the accuracy of LLMs inference. Details can be found in Appendix A.3. The results show that after adding the noise or conducting the quantization on the weights, the LLMs end up learning the concepts at slower paces and deeper layers.
@@ 2 Related Work
Figure 2: The LLMs are trying to understand easy tasks and complex tasks. The more complex the task, the deeper layers an LLM needs to understand. The stronger the LLM is, the more complex the task level it can learn.
@@ 2.1 Concepts Representation in DNNs
Identifying similarities across various examples to form concepts, plays a crucial role in human decision-making Armstrong et al. (1983). Many studies have explained DNNs’ (Deep Neural Networks) decision-making based on a conceptual perspective, describing the global behavior of DNNs in a human-understandable way Su et al. (2022); Räz (2023); Ren et al. (2023); Deng et al. (2021). For example, Yeh et al. (2019) demonstrated that DNNs exhibit conceptual representations through the activation patterns observed in their hidden or output layers. Further, Räz (2023) indicates that DNNs learned not only conceptual representations of predicted categories but also indirect concepts that contribute to the prediction. A notable study reveals the existence of a representation bottleneck, highlighting a cognitive disparity between DNNs and humans. This phenomenon is characterized by DNNs’ tendency to grasp representations of concepts that are either too simple or overly complex, while they often struggle with acquiring representations of concepts of moderate complexity Deng et al. (2021). Motivated by previous work, our paper aims to cover concepts of different complexities to understand concepts within LLMs further.
@@ 2.2 Knowledge and Concepts in LLMs
The impressive performance of the LLMs in various downstream tasks (e.g. LLMs can predict factual statements about the world based on prompts Meng et al. (2022)) has led to a great discussion about whether these capabilities are ‘stochastic parrots’ or LLMs understands these concepts. Pioneeringly, Gurnee and Tegmark (2023) showed that LLMs internally store concepts like latitude, longitude, and time. Similarly, another work showed that the internal states of LLMs can detect the truth of a statement Azaria and Mitchell (2023); Su et al. (2024b)Geva et al. (2023) also came to similar conclusions by artificially blocking or “knocking out" specific parts of the LLMs to observe their effects on the inference process. These related studies show the existence of structures for understanding concepts within LLMs, motivating us to explore how concepts at various complexities are encoded within various depths of LLMs.
@@ 2.3 Explorations of Interpretability in LLMs
Many related studies have deconstructed the inner layers of LLMs from various perspectives to understand the mechanisms inside such models Zhao et al. (2024)Fan et al. (2024) computes stopping signals by evaluating key features to early stop the LLM inference and get the internal performance of the LLMs, concluding that not all layers are necessary. Through pruning the LLMs, Gromov et al. (2024) found that the parameters of some layers were not utilized correctly. Men et al. (2024) also shows a high level of redundancy in the LLMs’ architecture. Probes trained with logistic regression are a well-established method Alain and Bengio (2016) that has been applied in classifying the truthfulness of LLMs and has been validated in many studies Marks and Tegmark (2023); Azaria and Mitchell (2023); Li et al. (2024). The latest work detects different layers in the Llama series responding to facts or counterfactuals by probing techniques Ju et al. (2024). Inspired by these works, we propose Concept Depth to summarize these phenomena. Our work focuses on the Concept Depth that appears in the LLMs, analyzing it experimentally by training linear classifier probes, which makes our work different from others.
@@ 3 Analyzing Method
In this paper, we design a probing framework to understand how concepts at various levels are encoded within LLMs and investigate whether the internal representations are robust to concepts. For instance, Figure 1 demonstrates the representation project of the Counterfact dataset.
@@ 3.1 Linear Classifier Probing
Probe technology Alain and Bengio (2016) is a method for analyzing and evaluating the internal representations of a neural network by applying a specific probe task, typically a classification or regression task, to a particular layer of the model. This technique measures the layer’s ability to represent information for the given task, thereby revealing the features and information captured by different layers of the model. Our approach involves extracting the representations from each layer of the large model, training a binary classifier on these representations, and validating its accuracy.
For one specific task $w$ that contains $n$ questions, the hidden feature set in LLMs is $x\in\mathbb{R}^{n\times d_{model}}$, where $n$ denotes number of samples, and $x^{(i)}\in\mathbb{R}^{1\times d_{model}}$ represent the representation at a certain layer, where $d_{model}$ donate the dimension for the hidden layer representation. Binary label $y^{(i)}$ is set as 0 or 1. The objective function of such a binary Logistic regression classifier probes with L2 regularization can be written as:
$J(\theta)=-\frac{1}{n}\sum_{i=1}^{n}Cost(\sigma(x^{(i)}),y^{(i)})+\frac{\lambda}{2n}\sum_{j=1}^{m}\theta_{j}^{2}$ (1)
$\begin{array}[]{cc}&Cost(\sigma(x^{(i)}),y^{(i)})=y^{(i)}\log\left(\sigma(\theta^{T}x^{(i)})\right)\\ &+(1-y^{(i)})\log\left(1-\sigma(\theta^{T}x^{(i)})\right)\end{array}$ (2)
where $\theta$ is the parameter in this Logistic regression model, $\lambda$ is the regularization parameter. The linear model predicts LLM’s response to the test set, compared with the true label. This yields a quantification of how well LLMs understand the current depth. If the binary model gets good accuracy at a certain layer, that means the LLM can distinguish true or false in this layer.
@@ 4 Experimental Setting
Our experiments used nine datasets containing three aspects (emotion understanding, reasoning, and fact-checking). We categorized these nine datasets from easy to complex levels according to the performance of LLaMA3-8B-Instruct Dubey et al. (2024) on each dataset (see Section 4.2.1) to anchor the difficulty of the datasets. Specifically, the datasets in which the linear probes can obtain high classification accuracy at the initial or middle depth of the LLMs are categorized as easy levels. Other datasets where linear probes can only accurately classify at a deeper layer of the model or even fail to classify accurately are categorized as complex levels. The overall accuracy of these datasets on the LLaMA3-8B-Instruct model was consistent with the probe results and had a significant correlation. In Section 4.1, we introduce the LLMs used for experiments. The nine datasets are described in Section 4.2.
@@ 4.1 Models
In this paper, we employ three open-source model families: Gemma (2B, 7B) Team et al. (2024), LLaMA-2 (7B, 13B) Touvron et al. (2023), and Qwen (0.5B, 1.8B, 4B, 7B, and 14B) Bai et al. (2023) to support our analysis. Section 1 shows the number of layers in each model. We choose a linear classifier probe for the experiments during the probing analysis. The ratio of the training set to the test set is 8:2, following the usual approach of LLMs probing classifier Duan et al. (2024); Pal et al. (2023). We extract feature representations from the final layer in the transformer at each layer of LLMs(e.g. 14-th ’post_attention_layernorm’ in Llama2-7B (32 Layers in total)) as input to the probing classifier. The other series of models follow a similar processing pattern.
@@
Gemma-2B 18 Qwen-4B 40 LLaMA-7B 32
Qwen-0.5B 24 Gemma-7B 28 Qwen-14B 40
Qwen-1.8B 24 Qwen-7B 32 LLaMA-13B 40
@@ 4.2 Datasets
Table 4 presents the nine datasets we use, on Fact (Cities Marks and Tegmark (2023), CommonClaim Casper et al. (2023), Counterfact Meng et al. (2022)), Emotion (STSA Kim (2014), IMDb Maas et al. (2011), Sarcasm Misra and Arora (2023), HateEval Manolescu et al. (2019)), and Reasoning (StrategyQA Geva et al. (2021), Coinflip Wei et al. (2022b)) for our experiments. A detailed description of the dataset can be found in the Appendix A.2.
4.2.1 Anchoring Difficulties of Each Dataset
To ascertain the learning difficulty of each dataset, we have utilized the LLaMA3-8B-Instruct Dubey et al. (2024) model. Our approach involves testing each sample in the datasets as a binary classification problem via a prompting way. The model generates a response for each sample, from which we infer a judgment, categorizing it as either "Yes" or "No". By comparing these judgments with the actual labels, we compute the accuracy for each dataset.
Table 2 presents the results of this analysis. The dataset with the highest accuracy is deemed the easiest dataset to classify. Conversely, the dataset with the lowest accuracy is considered the most difficult to classify. This method quantitatively measures the learning difficulty associated with each dataset.
Table 2: Accuracy on nine datasets based on LLaMA3-8b-Instruct.
Dataset Accuracy
Coinflip 0.5080
Counterfact 0.5277
Common 0.5606
Sarcasm 0.6575
StrategyQA 0.7035
HateEval 0.7640
Cities 0.7687
STSA 0.9030
IMDb 0.9365
4.3 Metrics for Accuracy Variation
Definition 1 (Variation Rate) Given an LLM probing classifier $M=\{q,y,z,d\}$ ($q$, $y$, $z$, and $d$ are the input question, ground truth binary label, predicted label and total amount of layers, respectively), it has the accuracy $\alpha_{i}$ at i-th layer:
$\alpha_{i}=\frac{1}{|z|}*\sum_{k=1}^{|z|}{[y_{k}=z_{k}]},i\in\{0,1,2,...,d-1\}$
We denote the variation rate $\beta_{i}$ where
$\beta_{i}=\alpha_{i}/\alpha_{i-1},i\in\{1,2,...,d-1\}$
We introduce two metrics to capture variations in accuracy: the jumping point and the converging point and define them by the given definition of variation rate.
Definition 2 (Jumping point) We denote the jumping point
$J(M,D)=\min\{\frac{i}{d}\}$
$\ s.t.\ \beta_{i}>=1.1,i\in\{1,2,...,d-1\}$
where M and $D=(q,y)$ are the LLM classifier and the dataset.
When a noticeable boost in accuracy is observed, the jumping point signals the model’s recognition of a dataset’s critical patterns.
Definition 3 (Converging Point) We denote the converging point
$C(M,D)=\max\{\frac{i}{d}\}$
$s.t.|\beta_{i}-1|<0.03,i\in\{1,2,...,d-1\}$
where $M$ and $D=(q,y)$ are the LLM classifier and the dataset.
As the accuracy plateaus or starts declining, the converging point indicates the model’s learning saturation or peak learning capacity from the dataset. Analyzing these metrics offers deeper insight into the model’s learning dynamics and adaptability to various data types.
Figure 4: Analysis diagrams of Section 5.2. The converge proportion and peak accuracy of each model over the nine datasets. (a) shows the converged proportion over the datasets. (b) shows the peak accuracy over the datasets.